Ethanolamine, C2H7NO, is a clear, viscous, hygroscopic liquid with an ammonia-like odor. It is used to...

  1. Ethanolamine, C2H7NO, is a clear, viscous, hygroscopic liquid with an ammonia-like odor. It is used to remove CO2 and H2S from natural gas but it is also found in polishes, hair waving solutions, and in emulsifiers (often in pharmaceuticals). Overexposure to this compound can cause irritation to the eyes, skin and respiratory system as well as lethargy. If a 0.25M solution of this compound is found to be 125 % ionized, determine the Kb and the Ka for this substance.

  1. Calculate the pH of a solution made by dissolving 5 grams of sodium nitrite (NaNO2) into 150 ml of     0.092 M solution of nitrous acid if the Ka of nitrous acid is 4.0 x 10-4.
  1. All parts of the hemlock plant contain the alkaloid coniine. Coniine (C8H17N) is an alkaline liquid with a pKb of 1 and a density of 0.844 g/ml. Ingestion of the substance may cause weakness, drowsiness, nausea, vomiting, labored respiration, paralysis and death. If an attempted murderer poured 2.5 ml of coniine into someone’s 500 ml water bottle, what would the pH be for the resulting solution? If the lethal dose for mice of 7 mg/kg of body weight is similar to that for humans, would a 120 pound athlete die from drinking the contents of the water bottle?
Answer #1

Question 2.

According to the Henderson-Hasselbulch equation:

pH = pKa + Log[NaNO2/HNO2]

pKa of HNO2 = -Log(Ka) = -Log(4*10-4) = 3.398

i.e. pH = 3.398 + Log{(5 g/69 g.mol-1)/(0.15 L * 0.092 mol/L)}

i.e. pH = 3.398 + 0.72

Therefore, pH = 4.118